OpenStudy (anonymous):

Find the exact value Given: f(0)=sin0 0=60 degrees Problem: f(0/2)

OpenStudy (anonymous):

@jagr2713

jagr2713 (jagr2713):

i dont understand this let me ask them @iambatman @ganeshie8 @SolomonZelman @Directrix

OpenStudy (loser66):

Let me help you to make it clear: Find the exact value Given \(f(\theta) = sin \theta \) when \(\theta = 60^0\) find \(f(\dfrac{\theta}{2})\)